OpenStudy (anonymous):

the altitude of a triangle is 2cm shorter than its base. THe area is 17.5cm^2. Find the base of the triangle.

OpenStudy (***[isuru]***):



OpenStudy (***[isuru]***):

since altitude is 2cm shorter than base altitude (A) = B - 2 \[Area \ \ of \ \ \triangle = \frac{ 1 }{ 2 } \times Altitude \times base\]

OpenStudy (***[isuru]***):

Area = (1/2) AB but A = (B - 2) therefor Area = (1/2) B(B-2) 17.5 = (1/2)B(B-2) now solve this for B can u do it ... or should i keep going ?

OpenStudy (anonymous):

I can do it thank you!

OpenStudy (***[isuru]***):

u r welcome!